I messed up a recipe for a holiday punch by putting in too little sprite. Right now I have 2 quarts of the punch and it is 40% sprite. How much more pure sprite should I put into the punch so that the mixture is 45% sprite?
Answer by TimothyLamb(4379)

---
40*2 + 100x = 45(x + 2)
40*2 + 100x = 45x + 90
55x = 10
x = 10/55
x = 0.181818 quart
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php